Operations Report – Mar 25th

Crew 207 Operations Report 25 Mar 2019

SOL: 9

Name of person filing report: Tom Baldwin

Non-nominal systems:

Generator: Radiator hoses and alternator need replacement on the generator. A contractor has been arranged to make these repairs within a few days.

Generator

Generator (hours run): 12.5 hours

Generator turned on (1930 last PM), SOC (100%)

Generator was turned off at (0700 this AM), SOC (100%)

Diesel Reading – 50 %
Propane Reading – 60 %
Ethanol Free Gasoline – 2 gallons.

Water (auxiliary tank) – Not in use
Water (static tank) – 200 gallons (replenished by David)
Auxiliary to Static tank transfer– Not applicable

Gallons transferred: Not Applicable

Water in Green Hab – 142 gallons
Water (loft) – Static to Loft Pump used – yes
Water Meter: 142050.1 gallons

Toilet tank emptied: No

Deimos rover used: No, not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Not used

Hours: 85.1 hours

Beginning charge: 100 %

Ending charge: 100 %

Currently charging: Yes

Opportunity rover used: Not used

Hours:54.9hours

Beginning charge: 100 %

Ending charge: 95 %

Currently charging: Yes (now 100%)

Curiosity rover used: Used

Hours: 84.7 hours

Beginning charge: 100 %

Ending charge: 97 %

Currently charging: Yes (now 100%)

Notes on rovers: not used

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): N/A

Reason for use: Not used

Oil Added?: N/A
ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar : None.

CrewCar used and why, where? Not used

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: EVA suits normal, with all batteries greater than 13 volts,

Summary of Hab operations: Per recent electrical anomalies, checked SOC of electrical system.

06:00 SOC 100 %, generator on.

07:30 SOC 100%, generator off

10.50 SOC 91%

Summary of Green Hab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Questions, concerns and requests to Mission Support: None

Sol Summary Report – Mar 25th

Summary Title : Day #9 (Out of Sim)

Author : Julielynn Wong

Mission Status : All nominal

Sol Activity Summary:

COM continued data collection for the thoracentesis and bike-powered 3D printing studies.

COM + XO continued 3D printing items in Level 1 of the Hab.

ENG + GHO signed out their duties to XO + COM.

ENG + GHO conducted testing of the new drone controller, completed surveys, and submitted qualitative feedback to manufacturer.

XO transported ENG/Astronomer + GHO/Journalist to GJT this PM.

Look ahead plan : Continue 3D printing projects, drone flight mapping, drone controller testing, and data collection for bike-powered 3D printing studies.

Anomalies in work : None

Weather : Sunny

Crew physical status : Fine

EVA : Drone flight testing in PM.

Reports to be filed : Operations, GreenHab, Astronomy

Support Requested : None

Journalist Report – Mar 24th

Journalist Report March 24

Journalist: Diane Rothberg

We are operating as a reduced crew today, as we sadly said goodbye to our two capable, young scientists, Tiffany and Kevin.

We are out of sim, but there is still the Hab to be run, and Tom did his operations duties, I took care of the plants in the Greenhab, and Julielynn compiled scientific data and took care of other administrative loose ends.

Tom and I got to take advantage of this glorious, sunny day by going for a hike. It was a real treat, being back in the “earth’s atmosphere”, and outdoors without the burden of a heavy space suit. The terrain here is awe-inspiring, oddly shaped hills and rock formations in every brown-red hue imaginable, and fossils from thousands of years ago everywhere!

We head out of here tomorrow, but what a privilege it has been to participate in this rare, and wonderful “Mars” experience. Thank you MDRS, and thank you, Medical Makers!

Operations Report – Mar 24th

Crew 207 Operations Report 24 Mar 2019

SOL: 8

Name of person filing report: Tom Baldwin

Non-nominal systems: Generator: Radiator hoses and alternator need replacement on the generator. A contractor has been arranged to make these repairs within a few days.

Generator

Generator (hours run): 11 hours

Generator was turned off at 8am, SOC 67 %

Diesel Reading – 50 %

Propane Reading – 61 %

Ethanol Free Gasoline – 2 gallons.

Water (auxiliary tank) – Not in use

Water (static tank) – 210 gallons (replenished by David)

Auxiliary to Static tank transfer– Not applicable

Gallons transferred: Not Applicable

Water in Green Hab – 142 gallons

Water (loft) – Static to Loft Pump used – yes
Water Meter: 142050.1 gallons

Toilet tank emptied: No

Deimos rover used: No, not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Not used

Hours: 85.1 hours

Beginning charge: 100 %

Ending charge: 100 %

Currently charging: Yes

Opportunity rover used: Not used

Hours:54.9hours

Beginning charge: 100 %

Ending charge: 95 %

Currently charging: Yes (now 100%)

Curiosity rover used: Used

Hours: 84.7 hours

Beginning charge: 100 %

Ending charge: 97 %

Currently charging: Yes (now 100%)

Notes on rovers: not used

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): N/A

Reason for use: Not used

Oil Added?: N/A

ATV Fuel Used: N7A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar : None.

CrewCar used and why, where? Not used

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: EVA suits normal, with all batteries greater than 13 volts,

Summary of Hab operations: Checked SOC of electrical system.

09:00 66%, generator off. Reported to David.

10:20 63%, generator off. Reported to David.

14:30 55% generator on. Reported to David.

15:20 Power failure. Reported to David

15:35 Power restored

15:55 55%, generator off.

16:55 48%, generator off

17:20 44% generator on

19:00 46% generator on

Summary of Green Hab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Questions, concerns and requests to Mission Support: None.

GreenHab Report – Mar 24th

Crew 207 GreenHab Report – 24 Mar 2019

GreenHab Officer: Diane Rothberg

Environmental control: Heat and cooling w/ambient air 10:00 hr

Shade Cloth (40%)

Average temperatures: RH

Low: 16.0 deg C 20%

High: 28 deg C 45%

Hours of supplemental light: 0 Hours

Daily water usage for crops: 6 gal

Water in Blue Tank: 142 gallons

Crops watered at 08:00 hours (3 gal) and 17:00 hours (3 gal)

TIME NOTED TEMPERATURE (DEG C) / HUMIDITY (%)

1) 8:00 AM 18 deg C / 34%

2) 10:35 AM 30 deg C / 26%

3) 14:10 PM 24 deg C / 20%

4) 17:00 PM 23 deg C / 20%

Changes to crops:

Narrative: None

Harvest: Basil: 10 gm

Support/ Supplies needed: none

GreenHab Report – March 24th

Crew 207 GreenHab Report – 24 Mar 2019

GreenHab Officer: Diane Rothberg

Environmental control: Heat and cooling w/ambient air 10:00 hr

Shade Cloth (40%)

Average temperatures: RH

Low: 16.0 deg C 20%

High: 28 deg C 45%

Hours of supplemental light: 0 Hours

Daily water usage for crops: 6 gal

Water in Blue Tank: 142 gallons

Crops watered at 08:00 hours (3 gal) and 17:00 hours (3 gal)

TIME NOTED TEMPERATURE (DEG C) / HUMIDITY (%)

1) 8:00 AM 18 deg C / 34%

2) 10:35 AM 30 deg C / 26%

3) 14:10 PM 24 deg C / 20%

4) 17:00 PM 23 deg C / 20%

Changes to crops:

Narrative: None

Harvest: Basil: 10 gm

Support/ Supplies needed: none.

Sol Summary Report – Mar 24th

Summary Title : Day #8 (Out of Sim)

Author : Julielynn Wong

Mission Status : All nominal

Sol Activity Summary:

Last night, Crew 207 successfully completed our check-out inspection with David. COM continued data collection for the thoracentesis study.

This AM XO transported HSO + SCI to GJT. ENG + GHO hiked up to the North Ridge to observe fossils.

COM continued 3D printing items in Level 1 of the Hab.

Look ahead plan : Continue drone flight mapping + data collection for thoracentesis and bike-powered 3D printing studies. COM to begin inputting thoracentesis study data values for remote analysis and manuscript preparation. XO to transport ENG + GHO to GJT in PM on Sol 9.

Anomalies in work : None

Weather : Sunny

Crew physical status : Fine

EVA : Drone mapping flights in AM out of sim.

Reports to be filed : Operations, GreenHab

Support Requested : None

Science Report – March 23rd

Crew 207

Research Report

Submitted by Crew Scientist Tiffany Ni

3D Printing Projects

Today is the last full day of MDRS for myself and our crew HSO. Our 3D-printing experiments continue to progress. We have completed printing all of our labware, International Space Station medical inventory items, STEM teaching aids, and are now printing some things here and there around MDRS, as well as other miscellaneous 3D printed projects.

To date, we have printed:

1. Western blot combs (18 well combs for making gels)
2. Microscope slide holder (reusable holder for square coverslips)
3. Sieve mesh (2.0mm holes)
4. Sieve structure (for holding the mesh net)
5. Mobile phone microscope clip
6. Sample box (for 96 tubes)
7. Stackable freezer bin
8. Salt spray chamber drain
9. Triangular straws (in a rainbow of colors)
10. Sieve mesh (5.0mm holes)
11. Sieve structure (for holding the mesh net)
12. Slide box holder (for storing microscope glass slides)
13. Test tube rack
14. Test tube holder
15. Sample rotator components
16. Small test tube holder (for the freezer)
17. Heat gun nozzle mount
18. Tongue depressors (in a rainbow of colors)
19. Stoma bag clip (used as a chip bag clip)
20. Full set of fraction tools (mathematics educational tools)
21. Funnels (small, medium, large)
22. Construction blocks
23. 12 tube separator (for Eppendorf tubes)
24. The side part of thoracentesis trainer
25. Customized finger splint (originally printed on the ISS0

During our time at MDRS, our 3D printed straws proved incredibly useful for our crew members. It has helped greatly in our attempts to reduce our water consumption. At MDRS, we have also demonstrated that a number of our 3D printed labware items are functional – as validated by two basic science research students.

Even though we will be leaving tomorrow morning, we will continue to stay in contact with our MDRS crew members and send 3D printable files for incoming crew members to print! (Perhaps some reusable bubble tea straws?)

Signing off,
Tiffany Ni
Crew 207 Scientist

Copyright © The Mars Society. All rights reserved. | Main Site